# Counterspeech the ultimate shield! Multi-Conditioned Counterspeech Generation through Attributed Prefix Learning

### v2

@@ Abstract
Counterspeech has proven to be a powerful tool to combat hate speech online. Previous studies have focused on generating counterspeech conditioned only on specific intents  [^1] While ‘intent’ has an established meaning in Natural Language Generation, we retain this term to maintain consistency with the IntentCONANv2. (single attributed). However, a holistic approach considering multiple attributes simultaneously can yield more nuanced and effective responses. Here, we introduce HiPPrO, Hierarchical Prefix learning with Preference Optimization, a novel two-stage framework that utilizes the effectiveness of attribute-specific prefix embedding spaces hierarchically optimized during the counterspeech generation process in the first phase. Thereafter, we incorporate both reference and reward-free preference optimization to generate more constructive counterspeech. Furthermore, we extend IntentCONANv2 by annotating all $13,973$ counterspeech instances with emotion labels by five annotators. HiPPrO leverages hierarchical prefix optimization to integrate these dual attributes effectively. An extensive evaluation demonstrates that HiPPrO achieves a $\sim 38\%$ improvement in intent conformity and a $\sim 3\%$, $\sim 2\%$, $\sim 3\%$ improvement in Rouge-$1$, Rouge-$2$, and Rouge-L, respectively, compared to several baseline models. Human evaluations further substantiate the superiority of our approach, highlighting the enhanced relevance and appropriateness of the generated counterspeech. This work underscores the potential of multi-attribute conditioning in advancing the efficacy of counterspeech generation systems.222 Warning: The materials presented in this paper might be disturbing or offensive.  [^3] Accepted in ACL 2025 Main Conference.
@@ 1 Introduction
The rise of the Internet has transformed social media platforms into hostile grounds for hateful comments targeting specific communities. Hate speech (HS) (Awal et al., 2021; Chakraborty and Masud, 2022) carries offensive statements that leverage stereotypes to spread harmful content. In contrast, counterspeech (CS) Benesch et al. (2016a); Wright et al. (2017) involves constructive responses to counteract hate speech by promoting positive dialogue, thus mitigating online hostility while preserving diverse perspectives Schieb and Preuss (2016); Chandrasekharan et al. (2017). As hateful comments surge Leetaru (2019); Masud et al. (2021, 2022), depending only on human-generated counterspeech becomes impractical. In this context, machine learning models appear essential for analyzing and generating counterspeech, offering a promising solution for automating the process. By leveraging this, content moderators can efficiently counter online hate Parker and Ruths (2023); Garg et al. (2023); Chung et al. (2023). Consequently, several interests Mathew et al. (2019); Qian et al. (2019); Fanton et al. (2021a); Bonaldi et al. (2022); Hengle et al. (2024) have intensified in the development of counterspeech generation methods.
The conceptualization of CS generation has evolved from a simple sequence-to-sequence problem to a more nuanced approach, acknowledging the diverse and context-dependent nature of hate speech Chung et al. (2019); Mathew et al. (2019); Sheng et al. (2020); Parker and Ruths (2023); Chung et al. (2023). This paradigm shift has fostered the development of advanced generative models incorporating stylistic and condition-guided elements, such as politeness, joyfulness, and detoxification, to create more effective counter-narratives Saha et al. (2022); Sheng et al. (2020). Recent research has introduced intent-specific CS generation, where established strategies guide the generation process to combat hate speech Gupta et al. (2023); Hengle et al. (2024); Benesch et al. (2016b).
@@
A significant portion of online hate speech consists of short, abusive statements Benesch et al. (2016a), and CS has shown potential in effectively countering such harmful content. While conventional approaches generate a single response per hate speech instance Zhu and Bhat (2021); Qian et al. (2019), recent studies highlight the benefits of tailoring CS to specific attributes for generating more diverse responses Gupta et al. (2023); Hengle et al. (2024). In practical scenarios, hateful comments often include multiple user intentions, necessitating the development of counterspeech that effectively addresses diverse attributes, resulting in more comprehensive and effective responses. For instance, single-attribute approaches typically produce one-dimensional responses by focusing solely on intent (e.g., being informative) while neglecting the emotional harmony required for persuasive communication. Real-world hate speech frequently needs a range of emotional responses, making it imperative to generate counterspeech that balances factual accuracy with emotional engagement.
@@
This study introduces an advanced pipeline for counterspeech generation, addressing the implicit nature of hate speech with responses aligned to multiple attributes. We focus on four primary intents – ‘positive’, ‘informative’, ‘questioning’, and ‘denouncing.’ and five emotion categories –‘anger’, ‘disgust’, ‘joy’, ‘sadness’, and ‘surprise’. Furthermore, we curate MultiCONAN, the largest collection of intent-emo-specific counterspeech, with $13,952$ responses countering $3,487$ instances of hate speech. We propose HiPPrO, a novel two-phase framework that first learns attribute-specific prefix embeddings (key-value pairs) and then applies preference tuning to generate constructive, non-toxic responses. Comprehensive evaluations using automated metrics and human evaluations demonstrate that HiPPrO consistently outperforms existing methods in CS generation across various criteria and achieves comparable performance with state-of-the-art LLMs like GPT-3.5 and GPT-4.444The source codes and datasets attached as supplementary and will be made public upon acceptance of the paper.
@@ 2 Related Work
The evolution of CS datasets has progressed from crowdsourced collections to expert-curated, intent-specific compilations Qian et al. (2019); Chung et al. (2019); Fanton et al. (2021a); Gupta et al. (2023); Hengle et al. (2024), while CS generation techniques have advanced from basic sequence-to-sequence models to sophisticated multi-phase pipelines with attribute control Zhu and Bhat (2021); Saha et al. (2019). These advancements have significantly improved the nuance and effectiveness of automated CS in promoting constructive dialogue Benesch et al. (2016c). Concurrently, parameter-efficient fine-tuning methods like Prefix Tuning Li and Liang (2021) and prompt tuning Lester et al. (2021) have emerged, modifying inputs while preserving language model parameters. Recent developments in preference tuning include RLHF’s application to instruction-following tasks Ouyang et al. (2022a), Direct Preference Optimization (DPO) Rafailov et al. (2023), and a unified approach omitting both reward and reference models Hong et al. (2024), addressing challenges in scalability and model sensitivity.
Counterspeech Intent and Emotion
@@ 3 The MultiCONAN Dataset
Due to the superior quality of counterspeech (see Appendix 8.3), we selected IntentCONANv2 Hengle et al. (2024) as the foundation for our work. IntentCONANv2, while valuable, has limitations that restrict its ability to fully capture the diversity and complexity of counterspeech. Its dependence on a single attribute, ‘intent,’ oversimplifies responses and fails to account for the emotional tone, which plays a critical role in shaping effective counterspeech. For instance, an informative counterspeech can convey vastly different emotional tones, such as joy or sadness, which are not captured in the original dataset. Building upon this, we introduce MultiCONAN, an enhanced version featuring additional emotion class labels for each counterspeech instance. MultiCONAN includes $13,973$ CS instances of IntentCONANv2, each tagged with one of five emotion classes: Anger (AN), Joy (JO), Disgust (DI), Sad (SA), and Surprise (SU) (See Table 1). The added emotion annotations enhance analysis granularity and facilitate the development of models that integrate both intent and emotional context. This annotation framework enables exploration of how emotional tone interacts with strategic intent and supports the creation of nuanced CS generation models Gupta et al. (2023); Hengle et al. (2024). MultiCONAN thus serves as a valuable resource for advancing research in CS generation, aiming to produce contextually appropriate and emotionally resonant responses. For detailed information on the annotation process and the statistics of the data set, refer to Appendices 8.1,8.2, and 8.4, respectively.
Figure 2: Our proposed model, HiPPrO, follows a two-phase pipeline. In Phase 1, we train the prefix parameters ($\alpha,\beta,\gamma,\delta$) associated with prefix adapters ($\mathcal{F_{\alpha}},\mathcal{F_{\beta}}$) on encoder side and ($\mathcal{F_{\gamma}},\mathcal{F_{\delta}}$) on the decoder side. For $\alpha$ and $\beta$, the model is trained with hate speech and intent, separated by </s> as input. While training $\gamma$ and $\delta$, the optimal parameters, $\alpha^{*}$ and $\beta^{*}$, are kept fixed, and it includes hate speech, intent, and emotion, separated by </s> as input $x$, this is also used as a prompt input during the second phase. In Phase 2, we apply preference tuning using an odd ratio loss, where the ground-truth counterspeech serves as the chosen candidate $y_{s}$, and the model-generated counterspeech is treated as the rejected candidate $y_{r}$.
@@ 4 Proposed Methodology
In this section, we elaborate on the inner workings and structural components of HiPPrO, a novel automated counterspeech generation framework. Here, we explain how our designed model can address the previous challenges by (i) generating multi-intentional conditioned counterspeech that can address hateful comments through semantic relevancy, and (ii) aligning it with qualitative human-generated responses through a reward and reference-model-free approach to ensure that the generation is contrastive (Figure 2).
@@
We use our curated MultiCONAN dataset for generating multi-intentional counterspeech generation. Considering our dataset as $\mathcal{D}=\{(h_{1},i_{1},e_{1},c_{1}),\ldots,(h_{n},i_{n},e_{n},c_{n})\}$, where $h_{i}\in\mathcal{H}$ is the $i$-th hate speech statement, $c_{i}\in\mathcal{C}$ is the counterspeech corresponding to $h_{i}$, and $i_{i}\in\mathcal{I}$ and $e_{i}\in\mathcal{E}$ are the intent and emotion categories of $c_{i}$, respectively. Our objective is to learn a stochastic counterspeech generation function $\psi:\mathcal{H}\times\mathcal{I}\times\mathcal{E}\rightarrow\mathcal{C}$, such that $c_{i}\sim\psi(\cdot|h_{i},i_{i},e_{i})$.
We address this problem by decomposing the counterspeech generation task into two phases. In the first phase, we focus on learning the prefix vectors for individual attributes and capturing their conditional dependencies using a two-step method. Initially, we learn the prefix vector for the intent category and collect the optimal prefix vectors. Subsequently, we add another prefix vector initialized with the previously learned values and optimized it for both the intent and emotion categories while keeping the model parameters from step one fixed. In the second phase, we employ a reward and reference model-free alignment method called the odds ratio preference optimization algorithm Hong et al. (2024). During this, we treat the counterspeech generated in phase one as the rejected data column and the actual ground truth counterspeech as the selected data column. For all our experiments, we utilize FLAN-T5 Chung et al. (2022) as the base model due to its robust reasoning and multi-task learning capabilities.
@@
In this section, we mathematically formalize the concept of hierarchical prefix learning. The prefix vectors, represented by tunable key-value pairs Vaswani et al. (2017), are introduced in two sub-phases. In the first sub-phase, we add $|V_{\mathcal{I}}|$ virtual prefixes with dimension $d$ across $l$ layers. The prefix adapters, $\mathcal{F_{\alpha}}$ and $\mathcal{F_{\beta}}$, introduce task-specific continuous vectors, $\alpha,\beta\in\mathbb{R}^{|V_{\mathcal{I}}|\times l\times 2d}$ to the encoder and decoder, respectively, guiding counterspeech generation according to user intent.
We maximize the expected log-likelihood and collect the optimal prefix adapters $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$ for the intent-guided counterspeech generation, where the hate speech and intent are sampled from $\mathcal{D}$, as denoted by,
@@
where $\pi^{ENC}_{\theta}$ and $\pi^{DEC}_{\theta}$ are the encoder and decoder part of the model, respectively. The input $X=[h;i;c]$, where $(h,i,c)\in\mathcal{D}$. We further add another set of adapters, $\mathcal{F_{\gamma}}$, and $\mathcal{F_{\delta}}$, on the encoder and decoder side, which add trainable prefix vectors $\gamma,\delta\in\mathbb{R}^{|V_{\mathcal{E}}|\times l\times 2d}$ on top of the previously trained $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$, where $|V_{\mathcal{E}}|$ is the number of virtual tokens for second sub-phase. These prefix vectors are responsible for guiding the counterspeech generation with intent and emotion-specific attributes. Here, the reformulated input is $X^{{}^{\prime}}=[h;i;e;c]$, where $(h,i,e,c)\in\mathcal{D}$. During training, we freeze the model parameters $\theta$ and the intent-specific prefix parameters $\alpha,\beta$ and maximize the expected log-likelihood to get optimal prefix adapters $\mathcal{F_{\gamma^{*}}},\mathcal{F_{\delta^{*}}}$ for optimal $\gamma^{*},\delta^{*}$ given by,
@@
Here, all $\alpha^{*},\beta^{*},\gamma^{*}$ and $\delta^{*}$ are responsible for the intent and emotion-conditioned counterspeech generation. Now, our model is ready to advance to Phase 2, where we apply a preference-tuning approach to all pre-trained prefix vectors. This step further optimizes the model to align counterspeech responses more closely with human-generated outputs.
@@
In this phase, we focus on optimizing our model to produce counterspeech instances that are both effective and non-toxic by employing a reward and reference-free preference tuning approach. Inspired by Hong et al. (2024), we use the odd ratio method to align the output with the ground truth preferences. This method self-penalizes the model output probabilities using its corresponding odds denoted as,
@@
where $\sigma$ denotes the sigmoid function. The final objective is to minimize $J_{final}$ by updating the continuous prefix vectors so that the model outputs can align more with the ground-truth counterspeech. For detailed information regarding computational resources and hyperparameter settings, please refer to Appendix 8.8 and Appendix 8.9, respectively. 555 The model was trained for up to 50 epochs using a callback, with a fixed training batch size and a learning rate of 4 and $1\times e^{-4}$, respectively.
Method Prompt/Adapter ROUGE (↑) M (↑) BS (↑) CoSim (↑) IC (↑) EC (↑) TC (↑) T (↓)
@@
GPT-3.5-Turbo ZS $0.239$ $0.052$ $0.160$ $0.249$ $0.868$ $0.585$ $0.533$ $0.783$ $0.978$ $0.0007$
GPT-3.5-Turbo FS $0.242$ $0.059$ $0.161$ $0.249$ $0.868$ $0.564$ $0.309$ $0.527$ $0.888$ $0.012$
GPT-4 ZS $0.221$ $0.037$ $0.145$ $0.229$ $0.864$ $0.583$ $0.545$ $0.790$ $0.976$ $0.030$
@@
HiPPrO ${}_{\mathrm{\textbf{VT=3}}}$ (Ours) PrefixTuning $0.273^{*}$ $0.082^{*}$ $0.199^{*}$ $0.242$ $0.879^{*}$ $0.567$ $0.929^{*}$ $0.706$ $0.897$ $0.087$

### v3

@@ Abstract
Counterspeech has proven to be a powerful tool to combat hate speech online. Previous studies have focused on generating counterspeech conditioned only on specific strategies (single attributed). However, a holistic approach considering multiple attributes simultaneously can yield more nuanced and effective responses. Here, we introduce HiPPrO, Hierarchical Prefix learning with Preference Optimization, a novel two-stage framework that utilizes the effectiveness of attribute-specific prefix embedding spaces hierarchically optimized during the counterspeech generation process in the first phase. Thereafter, we incorporate both reference and reward-free preference optimization to generate more constructive counterspeech. Furthermore, we extend IntentCONANv2 by annotating all $13,973$ counterspeech instances with emotion labels by five annotators. HiPPrO leverages hierarchical prefix optimization to integrate these dual attributes effectively. An extensive evaluation demonstrates that HiPPrO achieves a $\sim 38\%$ improvement in strategy conformity and a $\sim 3\%$, $\sim 2\%$, $\sim 3\%$ improvement in Rouge-$1$, Rouge-$2$, and Rouge-L, respectively, compared to several baseline models. Human evaluations further substantiate the superiority of our approach, highlighting the enhanced relevance and appropriateness of the generated counterspeech. This work underscores the potential of multi-attribute conditioning in advancing the efficacy of counterspeech generation systems.111 Warning: The materials presented in this paper might be disturbing or offensive. Our code is available on Github and dataset is open-sourced on Hugging-face.
@@ 1 Introduction
The rise of the Internet has transformed social media platforms into hostile grounds for hateful comments targeting specific communities. Hate speech (HS) (Awal et al., 2021; Chakraborty and Masud, 2022; Masud et al., 2021a, 2023) carries offensive statements that leverage stereotypes to spread harmful content. In contrast, counterspeech (CS) Benesch et al. (2016a); Wright et al. (2017); Singh et al. (2024) involves constructive responses to counteract hate speech by promoting positive dialogue, thus mitigating online hostility while preserving diverse perspectives Schieb and Preuss (2016); Chandrasekharan et al. (2017). As hateful comments surge Leetaru (2019); Masud et al. (2021b, 2022); Yadav et al. (2024), depending only on human-generated counterspeech becomes impractical. In this context, machine learning models appear essential for analyzing and generating counterspeech, offering a promising solution for automating the process. By leveraging this, content moderators can efficiently counter online hate Parker and Ruths (2023); Garg et al. (2023); Chung et al. (2023). Consequently, several interests Mathew et al. (2019); Qian et al. (2019); Fanton et al. (2021a); Bonaldi et al. (2022); Hengle et al. (2024) have intensified in the development of counterspeech generation methods.
The conceptualization of CS generation has evolved from a simple sequence-to-sequence problem to a more nuanced approach, acknowledging the diverse and context-dependent nature of hate speech Chung et al. (2019); Mathew et al. (2019); Sheng et al. (2020); Parker and Ruths (2023); Chung et al. (2023). This paradigm shift has fostered the development of advanced generative models incorporating stylistic and condition-guided elements, such as politeness, joyfulness, and detoxification, to create more effective counter-narratives Saha et al. (2022); Sheng et al. (2020). Recent research has introduced strategy-specific CS generation, where established strategies guide the generation process to combat hate speech Gupta et al. (2023); Hengle et al. (2024); Benesch et al. (2016b).
@@
A significant portion of online hate speech consists of short, abusive statements Benesch et al. (2016a), and CS has shown potential in effectively countering such harmful content. While conventional approaches generate a single response per hate speech instance Zhu and Bhat (2021); Qian et al. (2019), recent studies highlight the benefits of tailoring CS to specific attributes for generating more diverse responses Gupta et al. (2023); Hengle et al. (2024). In practical scenarios, hateful comments often include multiple user intentions, necessitating the development of counterspeech that effectively addresses diverse attributes, resulting in more comprehensive and effective responses. For instance, single-attribute approaches typically produce one-dimensional responses by focusing solely on strategy (e.g., being informative) while neglecting the emotional harmony required for persuasive communication. Real-world hate speech frequently needs a range of emotional responses, making it imperative to generate counterspeech that balances factual accuracy with emotional engagement.
@@
This study introduces an advanced pipeline for counterspeech generation, addressing the implicit nature of hate speech with responses aligned to multiple attributes. We focus on four primary strategies– ‘positive’, ‘informative’, ‘questioning’, and ‘denouncing.’ and five emotion categories –‘anger’, ‘disgust’, ‘joy’, ‘sadness’, and ‘surprise’. Furthermore, we curate MultiCONAN, the largest collection of strategy-emo-specific counterspeech, with $13,952$ responses countering $3,487$ instances of hate speech. We propose HiPPrO, a novel two-phase framework that first learns attribute-specific prefix embeddings (key-value pairs) and then applies preference tuning to generate constructive, non-toxic responses. Comprehensive evaluations using automated metrics and human evaluations demonstrate that HiPPrO consistently outperforms existing methods in CS generation across various criteria and achieves comparable performance with state-of-the-art LLMs like GPT-3.5 and GPT-4.
@@ 2 Related Work
The evolution of CS datasets has progressed from crowdsourced collections to expert-curated, strategy-specific compilations Qian et al. (2019); Chung et al. (2019); Fanton et al. (2021a); Gupta et al. (2023); Hengle et al. (2024), while CS generation techniques have advanced from basic sequence-to-sequence models to sophisticated multi-phase pipelines with attribute control Zhu and Bhat (2021); Saha et al. (2019). These advancements have significantly improved the nuance and effectiveness of automated CS in promoting constructive dialogue Benesch et al. (2016c). Concurrently, parameter-efficient fine-tuning methods like Prefix Tuning Li and Liang (2021) and prompt tuning Lester et al. (2021) have emerged, modifying inputs while preserving language model parameters. Recent developments in preference tuning include RLHF’s application to instruction-following tasks Ouyang et al. (2022a), Direct Preference Optimization (DPO) Rafailov et al. (2023), and a unified approach omitting both reward and reference models Hong et al. (2024), addressing challenges in scalability and model sensitivity.
Counterspeech strategy and Emotion
@@ 3 The MultiCONAN Dataset
Due to the superior quality of counterspeech (see Appendix 9.3), we selected IntentCONANv2 Hengle et al. (2024) as the foundation for our work. IntentCONANv2, while valuable, has limitations that restrict its ability to fully capture the diversity and complexity of counterspeech. Its dependence on a single attribute, ‘strategy,’ oversimplifies responses and fails to account for the emotional tone, which plays a critical role in shaping effective counterspeech. For instance, an informative counterspeech can convey vastly different emotional tones, such as joy or sadness, which are not captured in the original dataset. Building upon this, we introduce MultiCONAN, an enhanced version featuring additional emotion class labels for each counterspeech instance. MultiCONAN includes $13,973$ CS instances of IntentCONANv2, each tagged with one of five emotion classes: Anger (AN), Joy (JO), Disgust (DI), Sad (SA), and Surprise (SU) (See Table 1). The added emotion annotations enhance analysis granularity and facilitate the development of models that integrate both strategy and emotional context. This annotation framework enables exploration of how emotional tone interacts with strategic strategy and supports the creation of nuanced CS generation models Gupta et al. (2023); Hengle et al. (2024). MultiCONAN thus serves as a valuable resource for advancing research in CS generation, aiming to produce contextually appropriate and emotionally resonant responses. For detailed information on the annotation process and the statistics of the data set, refer to Appendices 9.1,9.2, and 9.4, respectively.
Figure 2: Our proposed model, HiPPrO, follows a two-phase pipeline. In Phase 1, we train the prefix parameters ($\alpha,\beta,\gamma,\delta$) associated with prefix adapters ($\mathcal{F_{\alpha}},\mathcal{F_{\beta}}$) on encoder side and ($\mathcal{F_{\gamma}},\mathcal{F_{\delta}}$) on the decoder side. For $\alpha$ and $\beta$, the model is trained with hate speech and strategy, separated by </s> as input. While training $\gamma$ and $\delta$, the optimal parameters, $\alpha^{*}$ and $\beta^{*}$, are kept fixed, and it includes hate speech, strategy, and emotion, separated by </s> as input $x$, this is also used as a prompt input during the second phase. In Phase 2, we apply preference tuning using an odd ratio loss, where the ground-truth counterspeech serves as the chosen candidate $y_{s}$, and the model-generated counterspeech is treated as the rejected candidate $y_{r}$.
@@ 4 Proposed Methodology
In this section, we elaborate on the inner workings and structural components of HiPPrO, a novel automated counterspeech generation framework. Here, we explain how our designed model can address the previous challenges by (i) generating multi-attribute conditioned counterspeech that can address hateful comments through semantic relevancy, and (ii) aligning it with qualitative human-generated responses through a reward and reference-model-free approach to ensure that the generation is contrastive (Figure 2).
@@
We use our curated MultiCONAN dataset for generating multi-attribute counterspeech generation. Considering our dataset as $\mathcal{D}=\{(h_{1},i_{1},e_{1},c_{1}),\ldots,(h_{n},i_{n},e_{n},c_{n})\}$, where $h_{i}\in\mathcal{H}$ is the $i$-th hate speech statement, $c_{i}\in\mathcal{C}$ is the counterspeech corresponding to $h_{i}$, and $i_{i}\in\mathcal{I}$ and $e_{i}\in\mathcal{E}$ are the strategy and emotion categories of $c_{i}$, respectively. Our objective is to learn a stochastic counterspeech generation function $\psi:\mathcal{H}\times\mathcal{I}\times\mathcal{E}\rightarrow\mathcal{C}$, such that $c_{i}\sim\psi(\cdot|h_{i},i_{i},e_{i})$.
We address this problem by decomposing the counterspeech generation task into two phases. In the first phase, we focus on learning the prefix vectors for individual attributes and capturing their conditional dependencies using a two-step method. Initially, we learn the prefix vector for the strategy category and collect the optimal prefix vectors. Subsequently, we add another prefix vector initialized with the previously learned values and optimized it for both the strategy and emotion categories while keeping the model parameters from step one fixed. In the second phase, we employ a reward and reference model-free alignment method called the odds ratio preference optimization algorithm Hong et al. (2024). During this, we treat the counterspeech generated in phase one as the rejected data column and the actual ground truth counterspeech as the selected data column. For all our experiments, we utilize FLAN-T5 Chung et al. (2022) as the base model due to its robust reasoning and multi-task learning capabilities.
@@
In this section, we mathematically formalize the concept of hierarchical prefix learning. The prefix vectors, represented by tunable key-value pairs Vaswani et al. (2017), are introduced in two sub-phases. In the first sub-phase, we add $|V_{\mathcal{I}}|$ virtual prefixes with dimension $d$ across $l$ layers. The prefix adapters, $\mathcal{F_{\alpha}}$ and $\mathcal{F_{\beta}}$, introduce task-specific continuous vectors, $\alpha,\beta\in\mathbb{R}^{|V_{\mathcal{I}}|\times l\times 2d}$ to the encoder and decoder, respectively, guiding counterspeech generation according to user strategy.
We maximize the expected log-likelihood and collect the optimal prefix adapters $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$ for the strategy-guided counterspeech generation, where the hate speech and strategy are sampled from $\mathcal{D}$, as denoted by,
@@
where $\pi^{ENC}_{\theta}$ and $\pi^{DEC}_{\theta}$ are the encoder and decoder part of the model, respectively. The input $X=[h;i;c]$, where $(h,i,c)\in\mathcal{D}$. We further add another set of adapters, $\mathcal{F_{\gamma}}$, and $\mathcal{F_{\delta}}$, on the encoder and decoder side, which add trainable prefix vectors $\gamma,\delta\in\mathbb{R}^{|V_{\mathcal{E}}|\times l\times 2d}$ on top of the previously trained $\mathcal{F_{\alpha^{*}}},\mathcal{F_{\beta^{*}}}$, where $|V_{\mathcal{E}}|$ is the number of virtual tokens for second sub-phase. These prefix vectors are responsible for guiding the counterspeech generation with strategy and emotion-specific attributes. Here, the reformulated input is $X^{{}^{\prime}}=[h;i;e;c]$, where $(h,i,e,c)\in\mathcal{D}$. During training, we freeze the model parameters $\theta$ and the strategy-specific prefix parameters $\alpha,\beta$ and maximize the expected log-likelihood to get optimal prefix adapters $\mathcal{F_{\gamma^{*}}},\mathcal{F_{\delta^{*}}}$ for optimal $\gamma^{*},\delta^{*}$ given by,
@@
Here, all $\alpha^{*},\beta^{*},\gamma^{*}$ and $\delta^{*}$ are responsible for the strategy and emotion-conditioned counterspeech generation. Now, our model is ready to advance to Phase 2, where we apply a preference-tuning approach to all pre-trained prefix vectors. This step further optimizes the model to align counterspeech responses more closely with human-generated outputs.
@@
In this phase, we focus on optimizing our model to produce counterspeech instances that are both effective and non-toxic by employing a reward and reference-free preference tuning approach. Inspired by Hong et al. (2024), we use the odd ratio method to align the output with the ground truth preferences. This method self-penalizes the model output probabilities using its corresponding odds, denoted as,
@@
where $\sigma$ denotes the sigmoid function. The final objective is to minimize $J_{final}$ by updating the continuous prefix vectors so that the model outputs can align more with the ground-truth counterspeech. For detailed information regarding computational resources and hyperparameter settings, please refer to Appendix 9.9 and Appendix 9.10, respectively. 222 The model was trained for up to 50 epochs using a callback, with a fixed training batch size and a learning rate of 4 and $1\times e^{-4}$, respectively.
Method Prompt/Adapter ROUGE (↑) M (↑) BS (↑) CoSim (↑) SC (↑) EC (↑) TC (↑) T (↓)
@@
GPT-3.5-Turbo ZS $0.239$ $0.052$ $0.160$ $\mathbf{0.249}$ $0.868$ $\mathbf{0.585}$ $0.533$ $0.783$ $0.978$ $\mathbf{0.0007}$
GPT-3.5-Turbo FS $0.242$ $0.059$ $0.161$ $\mathbf{0.249}$ $0.868$ $0.564$ $0.309$ $0.527$ $0.888$ $0.012$
GPT-4 ZS $0.221$ $0.037$ $0.145$ $0.229$ $0.864$ $0.583$ $0.545$ $\mathbf{0.790}$ $0.976$ $0.030$
@@
FLAN-T5XXL Retrieval-based $0.188$ $0.031$ $0.132$ $0.160$ $0.862$ $0.494$ $0.379$ $0.549$ $0.907$ $0.097$
GPT2XL Retrieval-based $0.113$ $0.011$ $0.074$ $0.154$ $0.821$ $0.366$ $0.309$ $0.441$ $0.754$ $0.103$
Llama-3.1-8B-Instruct Retrieval-based $0.128$ $0.026$ $0.086$ $0.197$ $0.828$ $0.462$ $0.347$ $0.678$ $0.891$ $0.096$
Mistral-7B-Instruct-v0.2 Retrieval-based $0.166$ $0.033$ $0.107$ $0.224$ $0.847$ $0.530$ $0.356$ $0.782$ $0.932$ $0.068$
DeepSeek-R1-Distill-Llama-8B Retrieval-based $0.132$ $0.027$ $0.086$ $0.210$ $0.830$ $0.471$ $0.373$ $0.751$ $0.890$ $0.059$
@@
DeepSeek-R1-Distill-Llama-8B PrefixTuning $0.035$ $0.0002$ $0.034$ $0.010$ $0.808$ $0.181$ $0.261$ $0.289$ $0.582$ $0.023$
DeepSeek-llm-7b-chat PrefixTuning $0.040$ 0.0010.001\ 0.0010.001 $0.035$ $0.027$ $0.805$ $0.144$ $0.261$ $0.311$ $0.484$ $0.019$
HiPPrO ${}_{\mathrm{\textbf{VT=3}}}$ (Ours) PrefixTuning $\mathbf{0.273^{*}}$ $\mathbf{0.082^{*}}$ $\mathbf{0.199^{*}}$ $0.242$ $\mathbf{0.879^{*}}$ $0.567$ $\mathbf{0.929^{*}}$ $0.706$ $0.897$ $0.087$